# Impact of Calibration and Position Errors on Astrophysical Parameters of the Hi 21cm Signal

### v1

@@ 9 Summary and Discussion
This study presents the development of an ANN and Bayesian-based framework for inferring astrophysical parameters from observed Hi power spectra (PS). We used the 21cm E2E pipeline to generate the mock observed power spectrum (PS) for different array configurations, including SKA Low core and HERA. However, generating a multitude of observable model signals across a multi-dimensional parameter space for Bayesian inference proved computationally difficult, especially with simulated observations. To address this, we adopted a formalism utilized by several groups, employing emulators for modelling the Epoch of Reionization (EoR) signal instead of actual simulations.
@@
For constructing the training datasets of observed PS, EoR 21cm signals were generated using the semi-numerical simulation 21cmFAST, with key EoR parameters (, , and ) described in section 2.1. These signals served as sky signals for mock observations using the 21cm E2E pipeline with various interferometric arrays like SKA Low and HERA, allowing us to construct sets of PS under different astrophysical conditions. These simulated observed PS were then used to train separate emulators for each interferometric array, facilitating Bayesian inference of the corresponding astrophysical parameters. A key advantage of these emulators, beyond conserving computational resources, is that our ANN-based models are trained directly on observed PS data while incorporating telescope effects. This allows for direct parameter inference without requiring separate corrections for telescope effects. In the first case of the study, we perform the inference using the developed ANN and Bayesian-based pipeline assuming perfect observing conditions where we assume the foreground had been perfectly removing only signal and telescope effect is there in the observed PS. We infer the astrophysical EoR parameters for the three sets for the signals constructed using the true astrophysical parameters, as listed in Tab.2. We have made the inference for the given different interferometers by considering the sample variance and the system noise at redshift = 9 corresponding to 1000 hrs observation time for each given interferometric array individually, detail result listed in Tab.2. Additionally, we have inferred astrophysical parameters from theoretical power spectra (PS) and compared these findings with parameters inferred from observed PS. In all three cases, the inferred parameters and closely matched the actual values, while the parameter was often not well-constrained, indicating high degeneracy. In the second case of the study, we assumed that the foregrounds were not perfectly removed, leaving some residual foreground due to calibration or position errors. This study aimed to determine the threshold level of calibration or position error for various interferometric arrays, such as SKA-Low and HERA. In a previous study by [1], 2D and 1D power spectra with gain and position errors were visually examined. They also calculated the root mean square (RMS) in the residual image across different percentage errors in calibration and position, comparing it with signal levels and thermal noise. They found the calibration error threshold to be 0.01 % and the position error threshold to be 0.48”.
@@
In this study, we delve deeper and systematically analyze these effects by performing inference on the observed power spectra from different interferometers. This allows us to understand the impact of various calibration and position errors on astrophysical parameters. The key findings noted from this study:
@@
Our analysis showed that the calibration error threshold for SKA-Low is 0.001%. Exceeding this threshold results in significant deviations in the inferred astrophysical parameters from their actual values, severely affecting the astrophysical processes. Similarly, for HERA, the calibration error threshold is also 0.001%, and surpassing this threshold leads to notable discrepancies between the inferred and actual parameters.
@@
The position error threshold for SKA-Low and HERA is found to be , as displacements cause significant deviations between the inferred astrophysical parameters and their actual values. This suggests that position errors can bias the inference of the astrophysical parameters.
@@
Among the two array layouts, we found that SKA Low exhibits slightly better tolerance than HERA for both calibration and position errors.
Thus, we conclude that for the existing interferometers, the calibration error should be nearly 0.001%, and the position error should be less than 0.48” to avoid biasing the inference. Beyond these thresholds, these errors start to affect the results. To reduce their impact, calibration must be performed with the same level of accuracy, or we need to develop modeling or mitigation techniques to eliminate the residual foregrounds arising from calibration or position errors, ensuring minimal influence on the inference. The simulations conducted in this study are based on several simplifying assumptions about foregrounds and instruments. The optimal tolerance for both calibration and position errors remains similar across both the scenarios. However, in the presence of calibration errors, SKA-Low’s power spectrum performance is slightly better than the HERA, with the inferred parameters more closely matching the actual values. This suggests the need for more detailed simulations, incorporating effects such as primary beam chromaticity, to better determine the optimal array configuration. Future studies will include these more complex simulations for a more comprehensive analysis.
@@ Acknowledgments
AT would like to thank the Indian Institute of Technology Indore for providing funding for this study in the form of a Teaching Assistantship. SM and AD acknowledge financial support through the project titled “Observing the Cosmic Dawn in Multicolour using Next Generation Telescopes” funded by the Science and Engineering Research Board (SERB), Department of Science and Technology, Government of India through the Core Research Grant No. CRG/2021/004025. The authors acknowledge the use of facilities procured through the funding via the Department of Science and Technology, Government of India sponsored DST-FIST grant no. SR/FST/PSII/2021/162 (C) awarded to the DAASE, IIT Indore.

### v2

@@ 9 Summary and Discussion
This study presents the development of an ANN and Bayesian-based framework for inferring astrophysical parameters from observed Hi power spectra (PS). We used the 21cm E2E pipeline to generate the mock observed power spectrum (PS) for SKA-Low core array configurations. However, generating a large number of modelled observable signals across a multi-dimensional parameter space for Bayesian inference is computationally expensive, particularly when using simulated observations. To address this, we adopted a formalism utilized by several groups, employing emulators for modelling the EoR signal instead of actual simulations.
For constructing the training datasets of observed PS, EoR 21cm signals were generated using the semi-numerical simulation 21cmFAST, with key EoR parameters (, , and ) described in section 2.1. These signals served as sky signals for mock observations using the 21cm E2E pipeline with interferometric arrays like SKA Low, allowing us to construct sets of PS under different astrophysical conditions. These simulated observed PS were then used to train emulator, facilitating Bayesian inference of the corresponding astrophysical parameters. A key advantage of our ANN-based emulators, beyond conserving computational resources, is their ability to perform direct parameter inference using observed power spectrum (PS) data while inherently accounting for telescope layout effects. In radio interferometric observations, limited and uneven uv-coverage introduces deviations from the theoretical PS through sampling artifacts, resolution loss, and convolution with the uv-sampling window. Traditionally, inferring astrophysical parameters from the observed PS requires applying separate corrections to account for these instrumental and sampling effects before comparing with theoretical models. In our approach, however, the ANN models are trained directly on observed-like PS data that already include all such instrumental and sampling-induced distortions. As a result, the emulator learns to incorporate these effects during training, allowing for parameter inference directly from the observed PS without requiring additional correction steps. This contrasts with methods relying on theoretical PS models, where observed data must first be adjusted to match idealized conditions to enable accurate parameter estimation.
@@
In the first case of the study, we perform the inference using the developed ANN and Bayesian-based pipeline assuming perfect observing conditions where we assume the foreground had been perfectly removing only signal and layout effect is there in the observed PS. We infer the astrophysical EoR parameters for the three sets for the signals constructed using the true astrophysical parameters, as listed in Tab.2. We have made the inference for the given different interferometers by considering the sample variance and the system noise at redshift = 9 corresponding to 1000 hrs observation time for each given interferometric array individually, detail result listed in Tab.2. Additionally, we have inferred astrophysical parameters from theoretical power spectra (PS) and compared these findings with parameters inferred from observed PS. In all three cases, the inferred mean values of the parameters and showed excellent agreement with the true input values, with the true values consistently falling within the 1 confidence intervals of the inferred posteriors (see Tab. 2). However, the parameter was often poorly constrained, indicating a high level of degeneracy. In the second case of the study, we assumed that the foregrounds were not perfectly removed, leaving some residual foreground due to gain calibration or sky model position errors. This study aimed to determine the threshold level of gain calibration or sky model position error for SKA-Low. In a previous study by [1], 2D and 1D power spectra with gain and position errors were visually examined. They also calculated the root mean square (RMS) in the residual image across different percentage errors in gain calibration and sky model position, comparing it with signal levels and thermal noise. They found the gain calibration error threshold to be 0.01 % and the sky model position error threshold to be 0.48”.
@@
In this study, we delve deeper and systematically analyze these effects by performing inference on the observed power spectra upcoming SKA-Low. This allows us to understand the impact of various gain calibration and sky model position errors on astrophysical parameters. The key findings noted from this study:
@@
Our analysis showed that the calibration error threshold for SKA-Low is 0.001%. Exceeding this threshold results in significant deviations in the inferred astrophysical parameters from their actual values, severely affecting the astrophysical processes. An equivalent calibration threshold was identified by [63].
@@
The position error threshold for SKA-Low is found to be , as displacements cause significant deviations between the inferred astrophysical parameters and their actual values. This suggests that sky model position errors can bias the inference of the astrophysical parameters.
@@
Thus, we conclude that for the existing interferometers, the gain calibration error should be nearly 0.001%, and the sky model position error should be less than 0.48” to avoid biasing the inference. Beyond these thresholds, these errors start to affect the results. To reduce their impact, calibration must be performed with the same level of accuracy, or we need to develop modeling or mitigation techniques to eliminate the residual foregrounds arising from calibration or sky model inaccuracies, ensuring minimal influence on the inference. It is important to note that the simulations presented in this study rely on several simplifying assumptions regarding foregrounds and instrumental effects. This underscores the need for more realistic and detailed simulations, including factors such as primary beam chromaticity and other instrumental systematics, to more accurately assess the robustness of parameter inference and to inform optimal array design. These complexities will be addressed in future work for a more comprehensive understanding.
@@ Acknowledgments
AT would like to thank the Indian Institute of Technology Indore for providing funding for this study in the form of a Teaching Assistantship. SM and AD acknowledge financial support through the project titled “Observing the Cosmic Dawn in Multicolour using Next Generation Telescopes” funded by the Science and Engineering Research Board (SERB), Department of Science and Technology, Government of India through the Core Research Grant No. CRG/2021/004025. The authors acknowledge the use of facilities procured through the funding via the Department of Science and Technology, Government of India sponsored DST-FIST grant no. SR/FST/PSII/2021/162 (C) awarded to the DAASE, IIT Indore. The authors would like to thank the anonymous reviewer and the scientific editor for their thoughtful comments and helpful suggestions, which have contributed to improving the quality of this work.